You have $150,000 to invest in a portfolio containing Stock X and Stock Y. Your goal is to create a portfolio that has an expected return of 10.35 percent. Stock X has an expected return of 9.54 percent and a beta of 1.24 and Stock Y has an expected return of 6.42 percent and a beta of .72.

  

How much money will you invest in Stock Y? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

What is the beta of your portfolio? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Investment in Stock Y = ___________ %

Portfolio Beta = __________ %
Answer #1
exp ret of Portfolio = Weight of Stock X*exp ret of Stock X+Weight of Stock Y*exp ret of Stock Y
10.35 = 9.54*Weight of Stock X+6.42*(1-weight of Stock X)
Weight of Stock X = 1.2596
Weight of Stock Y =1-weight of Stock X=1-1.2596=-0.2596

Amount invested in Stock Y =-0.2596153846*150000

=

-38942.31
Weight of Stock X = 1.2596
Weight of Stock Y = -0.2596
Beta of Portfolio = Weight of Stock X*Beta of Stock X+Weight of Stock Y*Beta of Stock Y
Beta of Portfolio = 1.24*1.2596+0.72*-0.2596
Beta of Portfolio = 1.38